What is the third one of 4 consecutive even integers , if the sum of these integers equals 92?
Answer by CubeyThePenguin(3113)

consecutive even integers: x, (x+2), (x+4), (x+6)

x + (x+2) + (x+4) + (x+6) = 92
4x + 12 = 92
4x = 80
x = 20

third integer = x + 4 = 24